OpenStudy (anonymous):

(cosx + 2sinx)^2 + (2cosx - sinx)^2=5

OpenStudy (anonymous):

what are the instructions?

OpenStudy (anonymous):

this about trigonometric identities and I do not know how to do the problem this

OpenStudy (whpalmer4):

\[(\cos x + 2\sin x)^2 + (2 \cos x- \sin x)^2 = 5\]\[(\cos x + 2\sin x)(\cos x + 2\sin x) + (2\cos x - \sin x)(2\cos x-\sin x) = 5\]Multiply that out and collect like terms. Remember that \[\sin^2 x + \cos^2 x = 1\]